lawrencewahkeleh:

Sheila is looking at some information for the obstacle course she is interested in completing. The x-coordinate is the number of the obstacle, while the y-coordinate is the average time to complete the obstacle, measured in minutes. (1, 8.25), (2, 9.075), (3, 9.9825), (4, 10.98075) Help Sheila use an explicit formula to find the average time she will need for the 8th obstacle

SamAlbert:

does this question have answer choices?

Mercury:

find an equation in y = mx + b form you can take two points (let's just use (1, 8.25), (2, 9.075) ) and apply the slope formula to find m then once you have m you can take one of the points, plug into y = mx + b, and solve for b finally, when you have a complete equation in the y = mx + b form, you can simply plug in x = 8 and find the appropriate y-value